OpenStudy (thekingofcr):

which quadratic equation has roots of -1+4i and -1-4i

563blackghost (563blackghost):

Can you please put up which equations we need to identify with?

OpenStudy (thekingofcr):

one sec

OpenStudy (thekingofcr):

x^2+2x+2=0 2x^2+x+17=0 x^2+2x+17=0 2x^2+x+2=0

OpenStudy (thekingofcr):

ok @563blackghost

563blackghost (563blackghost):

Plug them in quadratic formula. \(\LARGE\bf{x=\frac{-b \pm \sqrt{b^{2}-4ac}}{2a}}\) Remember the format of standard form is \(\large\bf{ax^{2} + bx+c}\). So identify `a, b, and c` and plug in. ~~~~ \(\LARGE\bf{x^{2}+2x+2=0}\) `a = 1 (due to only x)`, `b = 2` and `c = 2`. Plug. \(\LARGE\bf{x= \frac{-2 \pm \sqrt{2^{2}-4 \times 1 \times 2}}{2(1)}}\)

563blackghost (563blackghost):

This format applies to all equations.

OpenStudy (thekingofcr):

answer?

563blackghost (563blackghost):

You need to simplify with me. I can not give answers. It is against conduct. The equation above simplifies to... \(\LARGE\bf{x= \frac{-2 \pm 2i}{2}}\) In which we divide. \(\huge\bf{x={-1 \pm i}}\)

OpenStudy (eliesaab):

See the step by step solution given by Wolfram A;pha

563blackghost (563blackghost):

Or that.